MNO Co. issues zero coupon bonds on the market at a price of $334 per bond. Each bond has a face value of $1,000 payable at maturity in 17 years. The bonds are callable in 10 years at $683 . What is the yield to call for these bonds (in percent)? Answer to two decimals. Assume 1,000 par value and semi annual compounding
Answer #1

YTC = Couporadalur Pure forte YTC = - Coupon + 9 (CallValue-Purchase Price) noof years (CallValue+Purchase Price)

Since it is a zero coupon bond, no coupon, no of years = years to call i.e. 10 years. * 2 = 20 periods

0 (683-334) YTC = + 20 (683+334)

YTC = 3.43%

YTC annual = 3.43 * 2 = 6.86%  

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.